OpenStudy (anonymous):

Find dy/dx in terms of x & y. x^3+y^3=3xy^2 How do I even begin?

OpenStudy (astrophysics):

Implicit differentiation do you know it?

OpenStudy (astrophysics):

\[y' = \frac{ dy }{ dx }\]

OpenStudy (astrophysics):

So I'll start you off, \[3x^2 + 3y^2 \frac{ dy }{ dx } = ...\] now apply the product rule

OpenStudy (anonymous):

No, I did not know that. Thank you! Should I have 18xy + 9y^2 after applying the product rule?

OpenStudy (astrophysics):

Sorry my OS crashed

OpenStudy (astrophysics):

Applying the product rule on the left side we should have \[3y^2+6xy \frac{ dy }{ dx }\]

OpenStudy (astrophysics):

right side*

OpenStudy (astrophysics):

So we have \[3x^2+3y^2\frac{ dy }{ dx }=3y^2+6xy \frac{ dy }{ dx }\] now solve for dy/dx

OpenStudy (anonymous):

Oh okay that makes sense! Thank you so much!! I got -[(x)^2]/(2xy)

OpenStudy (astrophysics):

Mhm can you show me how you got that?

OpenStudy (anonymous):

\[\frac{ dy }{ dx } 3y ^{2} - 6xy = 3y ^{2}-3x ^{2}\] Then I moved the things from the left and created a fraction \[\frac{ 3y ^{2}-3x ^{2} }{ 3y ^{2}-6xy }\] After that, I simplified -3x^2/-6xy to -x^2/2xy

OpenStudy (astrophysics):

\[\frac{ dy }{ dx } = \frac{ x^2-y^2 }{ 2xy-y^2 }\] you get this yeah?

OpenStudy (astrophysics):

\[3x^2+3y^2\frac{ dy }{ dx }=3y^2+6xy \frac{ dy }{ dx } \implies 3x^2-3y^2 = 6xy \frac{ dy }{ dx }-3y^2\frac{ dy }{ dx }\]

OpenStudy (astrophysics):

\[x^2-y^2 = (2xy-y^2)\frac{ dy }{ dx } \implies \frac{ dy }{ dx } = \frac{ x^2-y^2 }{ 2xy-y^2 }\]

OpenStudy (anonymous):

I see what I did wrong. I assumed that you could just cancel the 3y^2/3y^2. I apologize and thank you so much for your time!

OpenStudy (astrophysics):

Yw :)